OpenStudy (anonymous):

Find the integral of (lnx^3)/x? Confused..

ganeshie8 (ganeshie8):

substitute \(u = \ln x \)

ganeshie8 (ganeshie8):

before that, simplify the log

ganeshie8 (ganeshie8):

use below property : \(\ln a^b = b\ln a\)

OpenStudy (anonymous):

ohhh okay that was what was making me stuck; the u! By the way.. if you're taking an integral of something and it's a product of two quantities, how do you take the integral?

ganeshie8 (ganeshie8):

there is no standard way for taking integral of product/quotient of two functions

OpenStudy (anonymous):

as opposed to the derivative being able to use the chain rule?

OpenStudy (anonymous):

oops sorry, product rule*

ganeshie8 (ganeshie8):

and u wont be successful most of the time

ganeshie8 (ganeshie8):

yes, there is not standard way for product of functions

ganeshie8 (ganeshie8):

we can take "derivative " of any complicated evil funciton manually.

ganeshie8 (ganeshie8):

but we cannot take "integral" of everything manually

ganeshie8 (ganeshie8):

and there are no product/quotient rules for integrals.

OpenStudy (anonymous):

got it :) thanks so much!

ganeshie8 (ganeshie8):

np :)